Q: Mary invested an amount of 20000 partly at the rate of 5% and partly at the rate of 8% interest per anum.If the amount of interest in the second case is 820$ more than the amount of interest earned in the first case,How much was invested at 5%?
Answer by mananth(16946)

Mary invested an amount of 20000 partly
x+y=20000
x at the rate of 5% and
y at the rate of 8% interest per annum.
If the amount of interest in the second case is 820$ more than the amount of interest earned in the first case,
8%y = 5%x+820
multiply by 100
8y=5x+82000
8y-5x=82000

x+y=100000---------------------(1)
8y-5x=82000-------------------(2)
13y=182000
y=14,000 at 8%

Multiply (1) by 5 and add to (2)
5x+5y=100000
8y-5x=82000
13y=182000
y=14,000
$14,000 invested at 5%?